OpenStudy (anonymous):

How do you convert 1.72 x 10^23 atoms of Potassium (K) to mass (in grams)?

OpenStudy (anonymous):

Using stoichiometry :) \[1.7\times10^{23} atoms \times \frac{ 1 mole K }{ 6.02\times10^{23} atoms}\times \frac{ 39.10 g K }{ 1 mole K }\] This will give you an answer. Stoichiometry is simple and very crucial in chemistry.